Hello,
Pls help me to solve this:

Select the pair of equation whose graphs are perpendicular;
A)2x - 8y = 9 & 12x - 3y = 7
B)3x + 6y = 8 & y = 2x - 8
C)y = 2x - 7 & x + y = 3
D)2y = -3x +5 & 2x + 3y = 4

Thanks
P.S:(This is the 2nd time that I am submitting this problem without receiving any answer)

Answer by nerdybill(7384)

Strategy is to find the slope of both equations. If the slopes are "negative reciprocal" the two lines are perpendicular. That is, the slopes multiplied results in -1.
.
A)2x - 8y = 9 & 12x - 3y = 7
slope of:
2x - 8y = 9
-8y = -2x+9
8y = 2x-9
y = (1/4)x-9/8 (slope is 1/4)
.
slope of:
12x - 3y = 7
-3y = -12x+7
3y = 12x-7
y = 4x-7/3 (slope is 4)
.
4(1/4) = 1 (NO, not perpendicular)
.
.
B)3x + 6y = 8 & y = 2x - 8
slope of:
3x + 6y = 8
6y = -3x+8
y = (-1/2)x+8/6 (slope is -1/2)
.
slope of:
y = 2x - 8 (slope is 2)
.
(-1/2)(2) = -1 (YES, it is perpendicular)
No need to solve the last two..
.
C)y = 2x - 7 & x + y = 3
D)2y = -3x +5 & 2x + 3y = 4